# multiGradICON: A Foundation Model for Multimodal Medical Image Registration

### v1

@@ 1 Introduction
Learning-based medical image registration [9, 53, 59, 5] has significantly improved over the last decade. Current SOTA learning-based deep registration networks [48, 47, 35, 36] are faster and more accurate than conventional registration methods using numerical optimization. However, learning-based approaches generally 1) focus on monomodal image registration and 2) are trained for a specific anatomical region (often the brain), making them much less generically applicable than conventional approaches. While many monomodal learning-based approaches have been developed, uniGradICON [47] is currently the only learning-based approach that is designed to be a universal method supporting registrations for different anatomies in one model working directly with images and SynthMorph [23] may generalize to other anatomies by training on synthetic shapes. Further, uniGradICON focuses on monomodal registration, and multimodal generalization is primarily achieved by instance optimization (IO). Hence, our goal is to further close the gap between conventional and learning-based registration approaches by developing multiGradICON, a multimodal generalization of uniGradICON. multiGradICON extends uniGradICON by 1) choosing a multimodal similarity measure, 2) incorporating multimodal registration tasks into training, and 3) exploring monomodal, multimodal, and randomized strategies for image similarity loss based on multimodal registration network inputs.
@@
    similarity loss randomization (i.e., randomly picking which modalities should be compared in the loss) improves registration accuracy for datasets containing multi-parametric data, even when scalar images are used for inference.
@@ 4 Results
Training design questions. We also investigate the impact of different factors such as training similarity loss selection ($1-\text{LNCC}$ or $1-\text{LNCC}^{2}$), instance optimization similarity loss selection ($1-\text{LNCC}^{2}$ or MIND-SSC), and training loss calculation strategy (baseline or label randomization). For this, we first train uniGradICON with $1-\text{LNCC}^{2}$ to investigate the effect of loss selection on generalization to multimodal pairs. Then, we introduce three variants of multiGradICON based on their image similarity loss calculation strategy: 1) the baseline multiGradICON-B approach which uses the same image pairing as input to the network and the loss; 2) multiGradICON-F which uses loss randomization but always samples from the same modality for the loss; 3) multiGradICON-R which also uses loss randomization but allows for sampling from different modalities. Finally, we obtain multiGradICON by further training our best-performing approach multiGradICON-R by including additional datasets to the training set (ThoraxCBCT, Pancreas, ACBD Diffusion (MD or FA)-Structure (T1w or T2w)). For this further training, we do not use $1-\text{CT}$ for data augmentation, we use lung-masked images for the COPDGene dataset, and we recompute the dataset weights (see Tab. 1). We report results for all our methods without instance optimization (w/o IO) or with 50 steps of IO using either $1-\text{LNCC}^{2}$ or MIND-SSC as similarity measures. Note that we perform all of the instance optimization operations on a given pair without any image similarity loss randomization, and we optimize the network parameters for the displacement field using an Adam optimizer with a learning rate of $2\times 10^{-5}$.
@@ 4.1 Performance on monomodal registration
multiGradICON - F 5.12 2.6e-4 5.56 2.0e-3 76.8 3.9e-5 40.5 5.1e-2 66.0 2.0e-2
multiGradICON - R 6.18 5.0e-4 6.60 3.6e-3 76.4 4.6e-5 40.2 2.7e-1 65.2 5.6e-2
multiGradICON 3.14 7.2e-4 5.85 4.0e-3 76.4 3.7e-5 39.5 8.6e-1 65.4 4.3e-2
1-LNCC2 uniGradICON 1.44 2.4e-4 2.80 1.3e-3 78.4 2.0e-4 52.9 9.4e-1 69.8 4.8e-2
@@
multiGradICON - F 1.69 1.4e-4 2.48 5.3e-4 78.4 4.9e-5 48.1 6.2e-1 69.3 1.8e-2
multiGradICON - R 1.78 5.9e-5 2.91 3.8e-4 78.1 6.7e-5 47.6 7.2e-1 68.4 3.6e-2
multiGradICON 1.63 1.2e-4 2.92 5.2e-4 78.2 7.6e-5 46.9 6.5e-1 68.2 3.7e-2
@@
multiGradICON - F 2.13 1.3e-5 3.39 5.5e-6 77.4 1.8e-7 44.5 8.6e-3 67.4 0
multiGradICON - R 2.25 0 3.92 7.4e-6 76.9 1.8e-7 44.4 3.6e-2 66.4 0
multiGradICON 2.03 1.3e-5 3.99 1.86e-6 77.1 7.4e-7 43.5 3.0e-2 66.4 0
@@
Datasets used for both uni- and multiGradICON training. The uni- and multiGradICON training datasets both contain lung (COPDGene CT), brain (HCP T1w MRI), abdomen (L2R Abdomen CT), and knee (OAI MRI) images. Tab. 2 shows that uniGradICON outperforms multiGradICON-B,F,R on the lung, abdomen, and knee datasets based on the initial prediction without instance optimization. This performance difference is around $\sim$3 mm for COPDGene, $\sim$8.5Dice score for the abdomen, and $\sim$3Dice score for the knee dataset. This result is expected, as uniGradICON is exclusively trained on these datasets and thus has better expertise in these areas. Conversely, multiGradICON is trained on diverse datasets where these specific tasks have lower weight during training. This is further supported by the brain registration results, where multiGradICON-B,F,R show similar performance, with multiGradICON-F even outperforming uniGradICON-$\text{LNCC}^{2}$ by a $\sim$0.2Dice score. Since brain datasets are more prevalent in the training sets (e.g., ABCD and BratsReg), multiGradICON performs similarly to uniGradICON on brain registration. After finetuning with anatomical region-based sampling, we observe a performance improvement on the COPDGene dataset, which forms 2.1% of the training set but is sampled at 8.3% during finetuning. The performance on the remaining datasets remains similar since their percentages do not change drastically. We observe similar performance improvement on the unseen NLST lung dataset (see Sec. 4).
@@
Instance optimization with $1-\text{LNCC}^{2}$ narrows the performance gap between uniGradICON and multiGradICON. The difference decreases to $\sim$0.19 mm for COPDGene, $\sim$6Dice score for the abdomen, and $\sim$1.6Dice score for the knee dataset. Instance optimization with MIND-SSC always under-performs instance optimization with $1-\text{LNCC}^{2}$ across all these datasets.
Different multimodal loss strategies perform differently for monomodal registration. Using the same modalities multiGradICON-F improves performance on the datasets to which this strategy is applied, compared to the baseline multiGradICON-B and random modality sampling in multiGradICON-R.
@@
Lung masking also affects registration performance. We train our multiGradICON variants using full lung CT images, whereas uniGradICON uses masked images that are zeroed out outside the lung. We observe that a registration model trained without lung masking cannot generalize to register fine details of the lung even if we provide masked lungs during inference. Therefore, during the finetuning process, we further train our model with region of interest (ROI)-masked lung images. After that, we achieve approximately 3 mm improvement in mTRE on masked lung registration without IO. Additionally, we observe approximately 0.75 mm performance improvement on non-masked lung images, even though we do not incorporate any non-masked lung images in the finetuning process. We hypothesize that both the sampling amount in diverse datasets and ROI-masked training are crucial for achieving good registration performance.
@@
multiGradICON - F 77.2 6.8e-5 2.94 7.4e-4 2.95 1.6e-3 2.73 9.2e-4 3.14 1.3e-3 45.5 1.8e-2 44.5 1.2e-2
multiGradICON - R 76.6 6.5e-5 3.07 1.2e-4 3.04 2.1e-3 2.87 1.3e-3 3.33 1.9e-3 43.6 4.4e-2 41.9 3.7e-2
multiGradICON 76.5 1.4e-4 3.06 8.8e-4 3.04 1.6e-3 2.90 1.2e-3 3.38 1.6e-3 43.8 2.3e-2 42.1 2.0e-2
1-LNCC2 uniGradICON 77.5 6.1e-4 2.93 7.7e-4 2.84 1.1e-3 2.48 9.4e-4 3.02 1.9e-3 47.0 8.5e-3 45.2 6.7e-3
@@
multiGradICON - F 77.6 1.3e-4 2.92 8.1e-4 2.80 1.2e-3 2.49 9.9e-4 2.95 1.6e-3 48.6 5.2e-3 47.6 6.7e-3
multiGradICON - R 77.2 1.6e-4 2.92 8.8e-4 2.79 1.1e-3 2.47 9.7e-4 2.99 1.6e-3 47.1 6.0e-3 45.4 8.1e-3
multiGradICON 77.2 2.6e-4 2.92 9.7e-4 2.81 1.2e-3 2.48 1.0e-3 2.99 1.9e-3 47.3 6.9e-3 45.8 1.0e-2
@@
multiGradICON - F 77.2 7.4e-7 2.72 0 2.52 0 2.23 0 2.63 3.9e-7 46.8 0 46.0 0
multiGradICON - R 76.6 0 2.69 0 2.53 1.3e-7 2.23 1.3e-7 2.64 6.2e-6 44.8 2.1e-7 43.5 0
multiGradICON 76.6 7.4e-7 2.69 0 2.53 0 2.22 2.6e-7 2.62 1.3e-7 45.1 1.6e-7 43.7 0
@@
multiGradICON - F 2.42 0 59.9 8.4e-3 71.6 2.9e-4 79.2 1.7e-3 80.5 2.3e-2
multiGradICON - R 2.66 0 58.9 3.8e-2 70.7 3.4e-4 78.5 2.3e-3 80.6 5.6e-2
multiGradICON 2.27 0 58.7 3.2e-3 71.0 1.8e-3 78.7 2.0e-3 81.8 2.1e-2
1-LNCC2 uniGradICON 1.77 8.7e-5 60.9 2.3e-1 70.4 1.5e-3 79.7 6.5e-3 82.2 2.4e-2
@@
multiGradICON - F 1.82 2.1e-4 61.4 4.3e-1 70.7 1.1e-3 79.9 6.9e-3 82.3 7.9e-3
multiGradICON - R 1.86 2.3e-4 60.2 2.7e-1 70.7 1.4e-3 79.6 7.1e-3 82.1 8.8e-3
multiGradICON 1.79 2.7e-5 60.6 2.8e-1 71.0 1.8e-3 79.6 6.3e-3 82.2 8.3e-3
@@
multiGradICON - F 1.95 0 63.5 1.4e-3 71.7 1.0e-6 79.1 0 81.8 0
multiGradICON - R 2.03 0 63.0 6.6e-6 71.6 3.2e-6 78.7 2.3e-6 81.7 2.3e-7
multiGradICON 1.94 0 63.4 0 71.8 4.3e-6 78.2 0 81.9 0
Unseen monomodal datasets. We also test on datasets that are never seen during training. Tab. 4 shows performance metrics for lung CT-CT NLST and brain T1w MRI registrations from the IXI and OASIS datasets. For NLST, uniGradICON achieves an mTRE of 2.07 mm, whereas the multiGradICON-B,F,R variants show a range between 2.42 and 2.74 mm. We observe a performance improvement on the NLST dataset of approximately 0.4 mm after finetuning.
Instance optimization with $1-\text{LNCC}^{2}$ decreases the performance gap to $\sim$0.02 mm on the NLST dataset. However, for brain registration, multiGradICON shows similar performance to uniGradICON, outperforming it by a $\sim$0.6Dice score on the IXI dataset and underperforming it by a $\sim$0.1Dice score on the OASIS dataset. These results show that multiGradICON scales well to monomodal tasks, providing performance close to that of uniGradICON, which specializes in monomodal registration.
@@ 4.2 Performance on multimodal registration
Multimodal training datasets. Our training dataset consists of several anatomical regions and modalities such as T1w, T1ce, T2w, FLAIR brain MRIs, CT abdominal scans, and fat and water-weighted DIXON images (see Tab. 1 for details). Tab.  5 shows registration performances for the HCP, Brats-Reg, Abdomen MR/CT, and UK Biobank datasets. In both the HCP and Brats-Reg datasets, we observe that uniGradICON fails to register pairs that contain images with large appearance differences. For instance, uniGradICON cannot register pairs containing T2w images, even with instance optimization. This is one of the key problems that we aim to solve with multiGradICON. We observe a significant performance improvement on the HCP ($\sim$59Dice score improvement), Brats-Reg, and MR-CT ($\sim$11Dice score improvement) datasets without instance optimization with our multiGradICON approach.
@@
multiGradICON - F 40.1 1.0e-3 9.10 2.2e-3 3.66 1.2e-3 5.06 1.4e-3 8.56 2.8e-3 6.50 3.2e-3 5.60 2.3e-3 58.2 6.0e-2 20.1 1.2e-2
multiGradICON - R 68.2 9.0e-4 3.58 2.3e-3 3.78 1.5e-3 4.48 2.1e-3 3.70 2.9e-3 4.90 3.8e-3 4.76 3.1e-3 61.1 8.4e-2 19.8 2.9e-2
multiGradICON 69.0 7.9e-4 3.64 2.5e-3 3.83 1.6e-3 4.68 1.9e-3 3.78 3.4e-3 5.05 4.1e-3 5.04 3.3e-3 61.8 4.0e-2 19.7 1.3e-2
1-LNCC2 uniGradICON 7.8 3.3e-3 12.06 3.2e-3 3.63 1.6e-3 5.13 2.0e-3 11.69 3.8e-3 6.81 4.8e-3 5.71 3.3e-3 75.1 6.9e-1 17.0 6.9e-3
@@
multiGradICON - F 71.9 4.9e-4 3.34 9.5e-4 3.55 1.5e-3 4.29 1.4e-3 3.39 1.5e-3 4.57 3.1e-3 4.78 2.6e-3 74.8 3.0e-1 18.9 6.9e-3
multiGradICON - R 72.8 5.7e-4 3.30 9.4e-4 3.55 1.5e-3 4.24 1.4e-3 3.37 1.4e-3 4.55 2.8e-3 4.70 2.5e-3 74.8 1.6e-1 18.6 6.6e-3
multiGradICON 73.1 9.3e-4 3.32 1.1e-3 3.56 1.8e-3 4.27 1.6e-3 3.38 1.8e-3 4.60 3.2e-3 4.72 3.0e-3 73.3 2.0e-1 18.6 8.0e-3
@@
multiGradICON - F 73.0 9.3e-7 2.72 0 2.62 0 2.87 6.7e-8 2.65 2.7e-7 2.94 2.0e-7 2.82 0 70.9 3.2e-4 25.2 0
multiGradICON - R 74.3 0 2.73 0 2.62 6.7e-8 2.88 6.7e-7 2.66 0 2.97 0 2.83 6.7e-8 70.3 1.0e-3 23.7 0
multiGradICON 74.4 1.4e-6 2.70 1.3e-7 2.62 1.2e-6 2.85 1.6e-6 2.62 0 2.94 1.3e-7 2.80 4.7e-7 70.8 2.7e-4 23.3 0
@@
We observe that multiGradICON-F cannot generalize to multimodal pairs compared to the other multiGradICON variants. It shows poor performance on several registration tasks such as T1w-T2w, T1ce-T2w, and T1ce-FLAIR brain registration. Although multiGradICON-F performs better than uniGradICON, these results suggest that sampling the same modality for similarity loss calculation causes a performance drop on multimodal datasets, while it improves the performance on monomodal pairs, as we discussed in Sec. 4.1. Additionally, we do not observe any significant multimodal performance improvement in uniGradICON when $1-\text{LNCC}^{2}$ is used as similarity loss during training. This indicates that using appropriate losses for multimodal registration may not lead to multimodal generalization if multimodal datasets were not used during training. We hypothesize that the diversity of the training dataset is more important for generalization.
@@
Unseen multimodal datasets. During the training of multiGradICON, we initially did not include the ThoraxCBCT and pancreatic CT-CBCT datasets in our training set. Although the performances of uniGradICON and multiGradICON-B,F,R on these datasets are close to each other, uniGradICON underperforms by $\sim$2Dice scores on the ThoraxCBCT dataset and outperforms by $\sim$0.5Dice score on the pancreatic CT-CBCT dataset. However, finetuning that incorporates the pancreatic dataset leads to a better Dice score for multiGradICON which outperforms uniGradICON by $\sim$0.7Dice score. The performance on the ThoraxCBCT dataset remains similar even when it is included in the training set during finetuning. We hypothesize that our model already converges on CT images during the initial training and that additional CBCT images do not affect the performance since they look sufficiently similar to CT images.

### v2

@@ 1 Introduction
Learning-based medical image registration [9, 53, 59, 5] has significantly improved over the last decade. Current SOTA learning-based deep registration networks [48, 47, 35, 36] are faster and more accurate than conventional registration methods using numerical optimization. However, learning-based approaches generally 1) focus on monomodal image registration and 2) are trained for a specific anatomical region (often the brain), making them much less generically applicable than conventional approaches. While many monomodal learning-based approaches have been developed, uniGradICON [47] is currently the only learning-based approach that is designed to be a universal method supporting registrations for different anatomies in one model working directly with images and SynthMorph [23] may generalize to other anatomies by training on synthetic shapes. Further, uniGradICON focuses on monomodal registration and multimodal generalization is primarily achieved by instance optimization (IO). Hence, our goal is to further close the gap between conventional and learning-based registration approaches by developing multiGradICON, a multimodal generalization of uniGradICON. multiGradICON extends uniGradICON by 1) choosing a multimodal similarity measure, 2) incorporating multimodal registration tasks into training, and 3) exploring monomodal, multimodal, and randomized strategies for image similarity loss based on multimodal registration network inputs.
@@
    similarity loss randomization (i.e., selecting random but identical modalities for image pairs for comparison in the loss) improves registration accuracy for datasets containing multi-parametric data, even when scalar images are used for inference.
@@ 4 Results
Training design questions. We also investigate the impact of different factors such as training similarity loss selection ($1-\text{LNCC}$ or $1-\text{LNCC}^{2}$), instance optimization similarity loss selection ($1-\text{LNCC}^{2}$ or MIND-SSC), and training loss calculation strategy (baseline or label randomization). For this, we first train uniGradICON with $1-\text{LNCC}^{2}$ to investigate the effect of loss selection on generalization to multimodal pairs. Then, we introduce three variants of multiGradICON based on their image similarity loss calculation strategy: 1) the baseline multiGradICON-B approach which uses the same image pairing as input to the network and the loss; 2) multiGradICON-F which uses loss randomization but always samples from the same modality for the loss; 3) multiGradICON-R which also uses loss randomization but allows for sampling from different modalities. Finally, we obtain multiGradICON by further training our best-performing approach multiGradICON-F by including additional datasets to the training set (ThoraxCBCT, Pancreas, ACBD Diffusion (MD or FA)-Structure (T1w or T2w)). For this further training, we do not use $1-\text{CT}$ for data augmentation, we use lung-masked images for the COPDGene dataset, and we recompute the dataset weights (see Tab. 1). We report results for all our methods without instance optimization (w/o IO) or with 50 steps of IO using either $1-\text{LNCC}^{2}$ or MIND-SSC as similarity measures. Note that we perform all of the instance optimization operations on a given pair without any image similarity loss randomization, and we optimize the network parameters for the displacement field using an Adam optimizer with a learning rate of $2\times 10^{-5}$.
@@ 4.1 Performance on monomodal registration
multiGradICON - F 5.63 2.6e-4 6.33 1.6e-4 75.2 1.3e-5 39.2 3.7e-2 65.3 2.0e-2
multiGradICON - R 5.76 4.5e-4 6.63 1.7e-4 74.9 4.8e-6 39.1 2.3e-2 65.8 7.7e-3
multiGradICON 3.29 1.0e-4 6.37 8.0e-4 76.3 1.0e-5 39.4 2.1e-2 66.5 5.8e-3
1-$\text{LNCC}^{2}$ uniGradICON 1.44 2.4e-4 2.80 1.3e-3 78.4 2.0e-4 52.9 9.4e-1 69.8 4.8e-2
@@
multiGradICON - F 1.73 1.1e-5 2.64 3.7e-5 78.1 1.9e-5 48.1 6.6e-1 69.3 1.3e-2
multiGradICON - R 1.76 7.4e-6 2.67 1.4e-4 77.8 4.0e-5 47.5 6.8e-1 68.9 1.1e-2
multiGradICON 1.63 3.7e-5 2.64 1.3e-4 78.2 2.8e-5 47.7 6.4e-1 69.4 1.1e-2
@@
multiGradICON - F 2.10 0 3.66 0 76.9 0 44.5 5.9e-3 67.2 0
multiGradICON - R 2.14 0 3.75 0 76.6 1.8e-7 44.7 8.5e-3 67.1 0
multiGradICON 1.95 0 3.70 0 77.2 0 44.7 5.4e-3 67.5 0
@@
Datasets used for both uni- and multiGradICON training. The uni- and multiGradICON training datasets both contain lung (COPDGene CT), brain (HCP T1w MRI), abdomen (L2R Abdomen CT), and knee (OAI MRI) images. Tab. 2 shows that uniGradICON outperforms multiGradICON-B,F,R on the lung, abdomen, and knee datasets based on the initial prediction without instance optimization. This performance difference is around $\sim$3 mm for COPDGene, $\sim$8.5Dice score for the abdomen, and $\sim$3Dice score for the knee dataset. This result is expected, as uniGradICON is exclusively trained on these datasets and thus has better expertise in these areas. Conversely, multiGradICON is trained on diverse datasets where these specific tasks have lower weight during training. This is further supported by the brain registration results, where multiGradICON-B,F,R show similar performance (within the range of  0.5 Dice score), with multiGradICON even outperforming uniGradICON by a $\sim$0.1Dice score. Since brain datasets are more prevalent in the training sets (e.g., ABCD and BratsReg), multiGradICON performs similarly to uniGradICON on brain registration. After finetuning with anatomical region-based sampling, we observe a performance improvement on the COPDGene dataset, which forms 2.1% of the training set but is sampled at 8.3% during finetuning. The performance on the remaining datasets remains similar since their percentages do not change drastically. We observe similar performance improvement on the unseen NLST lung dataset (see Sec. 4).
@@
Instance optimization with $1-\text{LNCC}^{2}$ narrows the performance gap between uniGradICON and multiGradICON. The difference decreases to $\sim$0.19 mm for COPDGene, $\sim$5Dice score for the abdomen, and $\sim$0.4Dice score for the knee dataset. Instance optimization with MIND-SSC always under-performs instance optimization with $1-\text{LNCC}^{2}$ across all these datasets.
Different multimodal loss strategies perform similarly for monomodal registration. There are no significant differences in performance, nor is there a clearly dominant method.
@@
Lung masking also affects registration performance. We train our multiGradICON variants using full lung CT images, whereas uniGradICON uses masked images that are zeroed out outside the lung. We observe that a registration model trained without lung masking cannot generalize to register fine details of the lung even if we provide masked lungs during inference. Therefore, during the finetuning process, we further train our model with region of interest (ROI)-masked lung images. After that, we achieve approximately 3 mm improvement in mTRE on masked lung registration without IO. We hypothesize that both the sampling amount in diverse datasets and ROI-masked training are crucial for achieving good registration performance.
@@
multiGradICON - F 76.1 3.3e-6 2.88 1.1e-4 2.82 3.4e-4 2.81 2.2e-4 3.00 2.1e-4 44.3 7.6e-3 43.6 2.0e-3
multiGradICON - R 75.9 7.4e-6 3.00 9.5e-5 2.92 4.0e-4 2.92 1.5e-4 3.15 1.5e-4 44.3 6.2e-3 43.4 9.4e-4
multiGradICON 76.5 7.2e-6 2.90 1.2e-4 2.86 3.6e-4 2.82 2.0e-4 3.05 2.0e-4 44.8 3.1e-3 44.1 7.6e-4
1-$\text{LNCC}^{2}$ uniGradICON 77.5 6.1e-4 2.93 7.7e-4 2.84 1.1e-3 2.48 9.4e-4 3.02 1.9e-3 47.0 8.5e-3 45.2 6.7e-3
@@
multiGradICON - F 77.3 4.6e-5 2.85 5.6e-4 2.74 1.1e-3 2.37 5.9e-4 2.91 1.0e-3 48.8 3.0e-3 48.2 3.2e-3
multiGradICON - R 77.1 3.5e-5 2.88 7.5e-4 2.75 1.2e-3 2.40 9.1e-4 2.91 1.4e-3 48.7 2.6e-3 48.0 2.9e-3
multiGradICON 77.3 7.7e-5 2.86 5.6e-4 2.75 1.0e-3 2.38 6.3e-4 2.91 1.0e-3 48.8 3.1e-3 48.1 3.9e-3
@@
multiGradICON - F 76.8 0 2.63 0 2.49 1.3e-7 2.18 0 2.57 0 46.6 0 46.2 0
multiGradICON - R 76.6 0 2.66 1.3e-6 2.52 0 2.22 1.3e-7 2.59 6.6e-7 46.4 0 45.8 0
multiGradICON 76.9 0 2.65 0 2.49 0 2.19 0 2.57 2.6e-7 46.7 0 46.2 0
@@
multiGradICON - F 2.98 0 57.1 5.0e-3 70.6 4.7e-5 78.2 8.3e-4 80.3 9.9e-3
multiGradICON - R 2.91 0 56.8 1.3e-3 70.3 1.9e-5 77.7 6.4e-4 80.4 4.7e-3
multiGradICON 2.25 0 58.0 2.5e-3 71.8 4.1e-5 78.4 8.4e-4 82.0 4.7e-3
1-$\text{LNCC}^{2}$ uniGradICON 1.77 8.7e-5 60.9 2.3e-1 70.4 1.5e-3 79.7 6.5e-3 82.2 2.4e-2
@@
multiGradICON - F 1.83 6.1e-4 60.2 2.7e-1 70.9 8.0e-4 79.3 5.8e-3 82.3 2.1e-3
multiGradICON - R 1.85 1.7e-5 60.6 3.1e-1 71.4 9.4e-4 79.3 6.6e-3 82.3 1.6e-3
multiGradICON 1.77 2.6e-4 60.6 2.9e-1 71.1 1.0e-3 79.4 5.2e-3 82.4 3.8e-3
@@
multiGradICON - F 2.01 0 64.1 0 72.6 1.2e-7 78.5 0 81.8 0
multiGradICON - R 2.03 0 63.0 0 72.7 0 78.3 3.1e-6 81.9 0
multiGradICON 1.93 0 63.9 0 72.9 8.9e-7 78.5 0 82.1 0
Unseen monomodal datasets. We also test on datasets that are never seen during training. Tab. 4 shows performance metrics for lung CT-CT NLST and brain T1w MRI registrations from the IXI and OASIS datasets. For NLST, uniGradICON achieves an mTRE of 2.07 mm, whereas the multiGradICON-B,F,R variants show a range between 2.74 and 2.98 mm. We observe a performance improvement on the NLST dataset of approximately 0.73 mm after finetuning.
Instance optimization with $1-\text{LNCC}^{2}$ closes the performance gap on the NLST dataset, with an mTRE of 1.77 mm. However, for brain registration, multiGradICON shows similar performance to uniGradICON, outperforming it by a $\sim$1.2Dice score on the IXI dataset and underperforming it by a $\sim$0.6Dice score on the OASIS dataset. These results show that multiGradICON scales well to monomodal tasks, providing performance close to that of uniGradICON, which specializes in monomodal registration.
@@ 4.2 Performance on multimodal registration
Multimodal training datasets. Our training dataset consists of several anatomical regions and modalities such as T1w, T1ce, T2w, FLAIR brain MRIs, CT abdominal scans, and fat and water-weighted DIXON images (see Tab. 1 for details). Tab.  5 shows registration performances for the HCP, Brats-Reg, Abdomen MR/CT, and UK Biobank datasets. In both the HCP and Brats-Reg datasets, we observe that uniGradICON fails to register pairs that contain images with large appearance differences. For instance, uniGradICON cannot register pairs containing T2w images, even with instance optimization. This is one of the key problems that we aim to solve with multiGradICON. We observe a significant performance improvement on the HCP ($\sim$59Dice score improvement), Brats-Reg, and MR-CT ($\sim$9.7Dice score improvement) datasets without instance optimization with our multiGradICON approach.
@@
multiGradICON - F 70.4 5.7e-6 3.04 1.5e-4 3.05 8.8e-5 3.46 1.3e-4 2.93 3.0e-4 3.49 2.9e-4 3.53 2.0e-4 59.4 9.4e-2 28.7 2.6e-3
multiGradICON - R 70.1 1.6e-5 3.19 1.2e-4 3.20 7.6e-5 3.57 7.8e-5 3.07 3.6e-4 3.63 4.1e-4 3.66 1.3e-4 61.8 2.2e-2 27.8 1.5e-3
multiGradICON 70.1 1.5e-5 3.18 1.4e-4 3.09 1.2e-4 3.52 1.3e-4 3.06 3.0e-4 3.60 3.4e-4 3.61 1.6e-4 59.7 3.3e-2 28.4 1.3e-03
1-$\text{LNCC}^{2}$ uniGradICON 7.8 3.3e-3 12.06 3.2e-3 3.63 1.6e-3 5.13 2.0e-3 11.69 3.8e-3 6.81 4.8e-3 5.71 3.3e-3 75.1 6.9e-1 17.0 6.9e-3
@@
multiGradICON - F 73.2 2.0e-4 3.14 8.6e-4 3.34 1.3e-3 3.90 9.3e-4 2.98 9.2e-4 4.35 1.5e-3 4.63 1.0e-3 75.8 4.4e-1 23.2 2.9e-3
multiGradICON - R 73.2 2.4e-4 3.15 1.0e-3 3.33 1.8e-3 3.87 1.4e-3 2.99 1.3e-3 4.36 2.2e-3 4.49 1.6e-3 76.6 3.0e-1 22.2 3.0e-3
multiGradICON 73.2 1.8e-4 3.17 1.0e-3 3.35 1.4e-3 3.89 1.0e-3 2.99 9.8e-4 4.40 1.8e-3 4.60 1.2e-3 75.5 4.4e-1 22.9 3.5e-3
@@
multiGradICON - F 75.1 0 2.60 0 2.48 0 2.71 2.6e-7 2.40 0 2.82 0 2.77 2.6e-7 69.5 1.4e-3 40.3 0
multiGradICON - R 75.0 0 2.65 0 2.51 5.3e-7 2.76 3.9e-7 2.42 0 2.85 1.3e-7 2.81 1.3e-7 71.9 5.6e-4 39.0 0
multiGradICON 75.1 0 2.60 0 2.51 1.3e-7 2.73 3.9e-7 2.41 0 2.82 2.6e-7 2.79 0 70.7 2.9e-3 40.1 0
@@
We observe that multiGradICON-F and multiGradICON-R outperform multiGradICON-B on the HCP and BraTS-Reg datasets, where loss randomization is applied. This demonstrates that loss randomization improves multimodal registration performance. Specifically, multiGradICON-F performs slightly better than multiGradICON-R across datasets. We hypothesize that selecting the same modality for loss calculations yields better results, because it simplifies the loss computations by only comparing within modalities, whereas comparing between different modalities would require a loss that can effectively work across modalities.
Additionally, we do not observe any significant multimodal performance improvement in uniGradICON when $1-\text{LNCC}^{2}$ is used as similarity loss during training. This indicates that using appropriate losses for multimodal registration may not lead to multimodal generalization if multimodal datasets were not used during training. We hypothesize that the diversity of the training dataset is more important for generalization.
@@
Unseen multimodal datasets. During the training of multiGradICON, we initially did not include the ThoraxCBCT and pancreatic CT-CBCT datasets in our training set. Although the performances of uniGradICON and multiGradICON-B,F,R on these datasets are close to each other, uniGradICON underperforms by $\sim$1Dice scores on the ThoraxCBCT dataset and outperforms by $\sim$0.5Dice score on the pancreatic CT-CBCT dataset. However, finetuning that incorporates the pancreatic dataset leads to a better Dice score for multiGradICON which outperforms uniGradICON by $\sim$0.9Dice score. The performance on the ThoraxCBCT dataset remains similar even when it is included in the training set during finetuning. We hypothesize that our model already converges on CT images during the initial training and that additional CBCT images do not affect the performance since they look sufficiently similar to CT images.